OpenStudy (anonymous):

3. You recently started the paperwork to purchase your new home, and were just notified that you can move into the house in 2 weeks. You decide to hire a moving company, but are unsure which company to choose. You search online and are interested in contacting two companies, Heavy Lifting and Quick Move, to discuss their rates. Heavy Lifting charges an $80 fee plus $35 per hour. Quick Move charges $55 per hour with no additional fees. a. Which mover provides a better deal for 2 hours of work? How did you arrive at your answer? b. Which mover provides a better deal for 15 hours of work? How d

OpenStudy (anonymous):

b. Which mover provides a better deal for 15 hours of work? How did you arrive at your answer? c. For what values h (hours) does Quick Move offer the better deal? Express your answer as an inequality. Explain how you reached your answer.

OpenStudy (anonymous):

35x+80 55x+0 a. 35(2)+80=150 55(2)=110 b. substitute 15 instead of 2 605 825 c. 35x+80=55x 80=20x x=4 at 4 hours they are equal. since u already found that at 2 hours quick move is better it would be: x<4

OpenStudy (anonymous):

oops, substitute h for where i used x

OpenStudy (anonymous):

b. I am confused about that one.

OpenStudy (anonymous):

see where i did (2)? i just multiplied by 15 instead

OpenStudy (anonymous):

so 35(15)+80=605 55(15)+0=825

OpenStudy (anonymous):

at 15 hrs the heavy lifting is better deal

OpenStudy (anonymous):

Thank you!!!!!

OpenStudy (anonymous):

i literally JUST covered this stuff in math class. :P

OpenStudy (anonymous):

lol. Yea I struggle with algebra. What school do you go to?

OpenStudy (anonymous):

Tidewater Community College

OpenStudy (anonymous):

oh cool. I am just glad to hear you are in college. LOL I have had some high schoolers on here helping me and they know it well and I feel a little slow. LOL

OpenStudy (anonymous):

well if it makes you feel better, I am 28 and been more than 8 years since last math. So I have to be taught at a very basic level. I feel so dumb.

OpenStudy (anonymous):

If you ever need it, I can use a voice server and can explain stuffs. Just give me a holler.

OpenStudy (anonymous):

LOL well I am 26 so don't feel bad.

OpenStudy (anonymous):

That would be awesome I can use all the help I can get and I am almost always on here.

OpenStudy (anonymous):

can i friend ya or whatever its called on facebook? would be easier for me

OpenStudy (anonymous):

or no, nm thats sounds creepy. cya around this site. :)

OpenStudy (anonymous):

lol you don't have to feel creepy as long as you aren't.

OpenStudy (anonymous):

I'm the type to walk a different way instead of behind a female, for fear she thinks I'm a stalker. Just a thing I do. :(

OpenStudy (anonymous):

well if you think fb is easier...that's fine...I mean I could always block you if I had to .LOL

OpenStudy (anonymous):

true... then yes please. nathan crain if you wanna search for me

OpenStudy (anonymous):

sent

OpenStudy (anonymous):

it's not t < 4 for part c you want all values of t such that 55t > 80 + 35t so 20t > 80 so t > 4

OpenStudy (anonymous):

then explain part a

OpenStudy (anonymous):

Quick Move was the better deal at 2 hrs. which is less than 4 hrs.

OpenStudy (anonymous):

oh, wait. i'm retarded. sorry. you want the cost to be SMALLER than the other company. you're right.

OpenStudy (anonymous):

I'm glad someone checked up on me though, tyvm for that.

OpenStudy (anonymous):

rachel, i think i saw you at Central Perk this one time

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Rachel, I used to live in ur town. I was at the army post. for like 3 years.

OpenStudy (anonymous):

rachel green is the name of friends character

OpenStudy (anonymous):

really!? Wow...small world huh?

OpenStudy (anonymous):

yea I know it is .

OpenStudy (anonymous):

extremely small world.

OpenStudy (anonymous):

how did you like it out here?